OpenStudy (anonymous):

the top walking deck of a sail boat is 15 3/4 feet long. this is only 5/8 of the length of the boat. how long is the boat????

OpenStudy (anonymous):

\[15 \frac{3}{4} \times \frac{8}{5}\]